# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

### v3

@@ Abstract
Knowledge-graph retrieval-augmented generation (KG-RAG) couples large language models (LLMs) with structured, verifiable knowledge graphs (KGs) to reduce hallucinations and expose reasoning traces. However, many KG-RAG systems compose multiple LLM modules (e.g planning, reasoning, and responding), inflating inference cost and binding behavior to a specific target KG. To address this, we introduce KG-R1, an agentic KG retrieval-augmented generation (KG-RAG) framework through reinforcement learning (RL). KG-R1 utilizes a single agent that interacts with KGs as its environment, learning to retrieve at each step and incorporating the retrieved information into its reasoning and generation. The process is optimized through end-to-end RL. In controlled experiments across Knowledge-Graph Question Answering(KGQA) benchmarks, our method demonstrates both efficiency and transferability: Using Qwen-2.5-3B, KG-R1 improves answer accuracy with fewer generation tokens than prior multi-module workflow methods that use larger foundation or fine-tuned models. Furthermore, KG-R1 enables plug and play: after training, it maintains strong accuracy on new KGs without modification. These properties make KG-R1 a promising KG-RAG framework for real-world deployment. Our code is publicly available at  https://github.com/Jinyeop3110/KG-R1.
@@ 1 Introduction
22footnotetext: Work done during an internship at MIT-IBM Watson AI Lab, IBM Research.
Figure 1: Overview of KG-R1, a multi-turn agentic framework for KG-RAG trained with reinforcement learning. The framework enables cost-efficient inference and demonstrates strong cross-KG transferability.
@@
Retrieval-augmented generation (RAG) has gained popularity as a way to enhance large language models (LLMs) with access to external knowledge, thereby reducing hallucinations and improving accuracy in knowledge-intensive tasks. Recent research has extended this idea to knowledge graph retrieval-augmented generation (KG-RAG), where knowledge graphs (KGs) are leveraged as the retrieval source. A KG is a structured representation of knowledge in the form of entities (nodes) and their relationships (edges) that encodes factual knowledge in a graph format. Augmenting LLMs with KGs has proven effective not only in mitigating the knowledge bottleneck, but also in improving reasoning performance over complex multi-hop relations and enhancing adaptation to continually evolving real-world information (Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c). These properties make KG-RAG especially promising in high-stakes domains, such as medical consultation and legal analysis (Xiong et al., 2023; Cui et al., 2024).
@@
As shown in Figure 1, typical KG-RAG adopts a modular workflow consisting of four primary subtasks: Retrieval to query facts from KGs, Reasoning to process the retrieved information, Reviewing to verify logical consistency, and Responding to synthesize the final answer (Baek et al., 2023; Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c). Each distinct subtask is handled by specialized LLM modules through two main methods: (i) prompt-based with task-specific instructions, often including in-context demonstrations (Baek et al., 2023); and (ii) fine-tuned modules tailored to particular tasks (e.g., SPARQL generation (D’Abramo et al., 2025; Lee & Shin, 2024) or relation extraction (Yao et al., 2019)) on specific KGs.
Figure 2: Prior multi-module methods are costly and do not transfer well across KGs. Left: mean end-to-end generated tokens per query on WebQSP (Yih et al., 2016). Right: average $\mathrm{Hit@1}$ over five out-of-training KGQA datasets (See Sec. 4.2). KG-R1 achieves both low token cost and strong cross-KG transferability.
@@
Despite improved reasoning accuracy, the real-world deployment of such workflows faces two key challenges, high computational cost and lack of generalization to new or updated KGs. First, prompt-based methods that repeatedly call large foundation LLMs accumulate inference across stages and drive up latency, token usage, and compute (e.g., ToG and ReKnoS; see left panel of Fig. 2). Second, these prompted or fine-tuned modules are typically tuned to a particular KG’s domain and schema (entity types, relations, constraints), often via curated in-context examples or KG-specific fine-tuning. As a result, performance does not transfer reliably when the domain shifts, the schema changes, or the system is deployed on a new KG (e.g., RoG; see right panel of Fig. 2).
@@
To tackle these challenges, we introduce KG-R1, an agentic KG-RAG system that employs end-to-end multi-turn reinforcement learning (Jin et al., 2025; Zeng et al., 2025; DeepSeek-AI, 2025). As shown in Figure 1, the architecture of KG-R1 has two components: a single LLM agent and a KG retrieval server (as an environment). The KG retrieval server hosts the knowledge graph along with a set of retrieval actions. The LLM agent iteratively performs cycles of short reasoning followed by retrieval actions over multiple turns, with each decision informed by knowledge obtained from the KG retrieval server, and generates a final answer. Figure 2 demonstrates that KG-R1 achieves both high efficiency and strong cross-KG transferability simultaneously using a 3B-parameter model, outperforming prior methods. Our key contributions are summarized as follows:
@@
1. KG-R1 framework. We introduce an agentic KG-RAG system (Section 3) that replaces multi-module pipelines with a single agent for KG-RAG, running against a lightweight KG server. The agent alternates between reasoning and retrieval actions over multiple turns, with the end-to-end trajectory optimized by RL using both turn-wise and outcome-based reward signals. Turn-wise rewards evaluate individual action effectiveness and adherence to formatting, while global rewards measure answer quality and retrieval relevance.
@@
2. Efficient inference. By consolidating reasoning and retrieval into a single-agent, near-continuous workflow, KG-R1 achieves competitive reasoning accuracy with a small-parameter model while reducing token usage. This lowers latency and computational cost, making deployment feasible under tight budgets. Experiments demonstrate improved performance and efficiency compared to traditional multi-module workflows (Section 4.1).
@@
3. Plug-and-play transferability across diverse KGs. KG-R1 easily transfers to diverse KGs and maintains strong KG-RAG performance (Section 4.2). The trained KG-R1 agent generalizes to new KGs without modification—backend KGs can be swapped without changing prompts or hyper-parameters, and without fine-tuning. This enables zero-shot transfer to unseen knowledge graphs.
@@ 2 Related Work
KG-RAG. Knowledge Graph Retrieval-Augmented Generation (KG-RAG) augments LLMs with structured knowledge graphs (KGs) to improve factuality and compositional reasoning. Early work grounds LLMs’ generation by translating natural-language questions into executable graph queries (e.g., SPARQL/Cypher), retrieving relevant subgraphs or answers, and feeding them back to the model (Ouyang et al., 2022; Izacard et al., 2023; Lee & Shin, 2024). More recent approaches adopt a modular LLM pipeline over KGs, interleaving natural-language reasoning with multi-stage planning, path search, and execution, where each stage uses a prompted or fine-tuned LLM (Luo et al., 2024; Sun et al., 2024; Wang et al., 2025c). Despite these advances, most systems rely on hand-engineered modules or prompt heuristics tied to a specific KG schema, which induce cost inefficiency and limit generalization. These challenges motivate our KG-R1 framework: a single-agent KG-RAG approach that improves efficiency and transfers well to new KGs.
Multi-turn RL for LLM. Reinforcement learning (RL) has become central to equipping LLMs with step-by-step (chain-of-thought) reasoning behavior (OpenAI et al., 2024; DeepSeek-AI, 2025). RL-enhanced models yield substantial gains in math and coding (Le et al., 2022; Chervonyi et al., 2025), and broader complex reasoning tasks. More recently, RL has been applied to agentic LLMs that invoke external tools (e.g., bash terminals and APIs) or interact with knowledge bases, improving tool use (Qin et al., 2024) and retrieval-augmented generation (RAG) (Jin et al., 2025; Wang et al., 2025a) to facilitate the tool use or RAG. Building on these advances, our work, KG-R1 adopts end-to-end RL as the core optimization algorithm for agentic KG-RAG framework.
@@ 3.1 Problem Definition
KG–RAG spans many applications, including conversational assistants (Chaudhuri et al., 2021), recommendation systems (Wang et al., 2025b), and open-domain QA (Zhu et al., 2025). In this work, we instantiate and evaluate our approach on Knowledge Graph Question Answering (KGQA), which provides a grounded testbed for KG-RAG: ground-truth answers are tied to a fixed KG, evaluation is verifiable and intermediate graph reasoning is interpretable. We now formalize knowledge graphs and KGQA tasks.
@@
Knowledge Graph Question Answering (KGQA) is a reasoning task based on KGs. Consider a dataset $D=\{(q,G,A_{q})\}$, where $q$ is a natural language question, $G$ is a KG, and $A_{q}\subseteq\mathcal{E}$ is the ground-truth answer set paired with the question. For each gold answer $e^{*}\in A_{q}$, there exist one or more reasoning paths in $G$ that connect anchor entities mentioned in $q$ to $e^{*}$. A reasoning path is a sequence of labeled edges $r_{1},\dots,r_{\ell}\in\mathcal{R}$ instantiated by entities $e_{0},\dots,e_{\ell}$ such that $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},$ where $e_{0}$ is an anchor entity mentioned in $q$ and $e_{\ell}=e^{*}\in A_{q}$, and $\mathcal{Z}(q,G,A_{q})$ denote the set of all valid reasoning paths for question $q$ over $G$ with respect to ground-truth answers $A_{q}$. In solving KGQA, given $q$ and $G$, a model attempts to discover a subset of valid reasoning paths from $\mathcal{Z}(q,G,A_{q})$ and predicts $\hat{A}_{q}$ based on the terminal entities of the discovered paths. The model’s performance is evaluated by comparing $\hat{A}_{q}$ with the ground-truth $A_{q}$. As an example, to answer the question “What is the capital of the U.S. state whose largest city is Chicago?”, a valid reasoning path is
$\text{Chicago}\xrightarrow{\texttt{located\_in\_state}}\text{Illinois}\xrightarrow{\texttt{capital}}\text{Springfield}$
@@ 3.2 KG-R1 Framework
Figure 3: KG-R1 framework: a single LLM agent undergoes multi-turn generation–execution loop with a schema-agnostic KG retrieval server and responds with the final answer.
KG-R1 casts KG-RAG as a multi-turn interaction with a KG interface (KG retrieval server). We prioritize two design principles. First, we design a single-agent architecture that simplifies deployment and enables efficient, low-cost inference. Second, we build a schema-agnostic KG retrieval server that avoids KG-specific assumptions and remains portable across varied KGs. Given a KGQA dataset $D=\{(q,G,A_{q})\}$, we set up a KG retrieval server and a KG-R1 agent.
KG retrieval server. The server hosts the knowledge graph $G$ and provides a set of retrieval actions $a\in\mathcal{U}_{\text{ret}}$ that enable graph traversal through 1-hop operations:
@@
The retrieval action set $\mathcal{U}_{\text{ret}}$ is sufficient for realizing any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ as a finite action sequence that arrives at $e_{\ell}\in A_{q}$ (i.e., the terminal entity is in the answer set). Forward traversal along $Z$ is implemented by $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ for $i=1,\dots,\ell$ and backward traversal is implemented by $\mathtt{get\_head\_entities}(r_{i},e_{i})$. Notably, this choice of the retrieval action set guarantees completeness and schema-free transferability, as formalized by propositions 3.1 and 3.2, whose proofs are provided in Appendix A.1.
Proposition 3.1 (Retrieval Action Set Completeness).
For any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ in $G$, there exists an action sequence in $\mathcal{U}_{\text{ret}}$ of length at most $\ell{+}1$ whose output includes $e_{\ell}$.
Proposition 3.2 (Schema-Free Transferability).
The semantics of $\mathcal{U}_{\text{ret}}$ depend only on directed triples $(e,r,e^{\prime})$, so any fixed client policy transfers across heterogeneous directed KGs by replacing $G$ without redefining $\mathcal{U}_{\text{ret}}$.
KG-R1 Agent. We model a single LLM agent that interacts with a KG retrieval server in a multi-turn setting. The agent undergoes initialization followed by a loop of generation and execution (see Figure 3). At initialization, a lightweight base LLM is configured with an instruction prompt $p$ (see the box below for the prompt template) containing general reasoning instructions, the user question $q$, and the KG retrieval server instructions (Table D.2). At each turn $t\leq H$, where $H$ is the maximum turn limit, the agent first undergoes the generation phase where it produces a response $m_{t}$ comprising two parts: (1) an internal reasoning wrapped in <think>...</think>, and (2) an action, which is either a KG retrieval action wrapped in <kg-query>...</kg-query>, or a final answer wrapped in <answer>...</answer>.
@@
In the following execution phase, we parse $m_{t}$ into an action $a_{t}\in\mathcal{U}_{\text{ret}}\cup\{\operatorname{Answer}\}$ using an exact-match parser $\Psi$ (i.e., $a_{t}=\Psi(m_{t})$). If $a_{t}\in\mathcal{U}_{\text{ret}}$, executing it yields an observation $o_{t+1}$ (i.e., retrieved entities, relations, or an error message in case the action does not generate a valid retrieval), which is appended to the dialogue context prior to the next turn. If $a_{t}=\operatorname{Answer}$, the interaction terminates: the content inside <answer>...</answer> is extracted and post-processed (normalization, deduplication, and entity resolution) to produce the predicted answer set $\hat{A}_{q}$. Given interaction trajectory $\tau=(p,\,(m_{1},o_{1}),\ldots,(m_{t},o_{t}))$ with $t\leq H$, the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ is defined over token sequences $m_{t}$ and governs how textual responses are generated.

### v5

@@ Abstract
Knowledge-graph retrieval-augmented generation (KG-RAG) couples large language models (LLMs) with structured, verifiable knowledge graphs (KGs) to reduce hallucination and provide reasoning traces. However, current KG-RAG systems often rely on fixed pipelines of multiple LLM modules (e.g., planning, reasoning, and responding), which inflate inference costs and tie performance to specific graph schemas. To address this, we introduce KG-R1, an agentic framework that optimizes KG-RAG through reinforcement learning (RL). Unlike modular workflows, KG-R1 uses a single agent that interacts with KGs as its environment, learning to retrieve information at each step and incorporating it into its reasoning and generation in a unified process. Across Knowledge-Graph Question Answering (KGQA) benchmarks, KG-R1 demonstrates both efficiency and transferability—using Qwen 2.5-3B, KG-R1 improves answer accuracy with fewer generation tokens than prior multi-module workflow methods that use much larger foundation or fine-tuned models. Furthermore, KG-R1 exhibits strong plug-and-play capability: after training, maintaining accuracy on unseen KGs without retraining. These properties make KG-R1 a promising KG-RAG framework for real-world deployment. Our code is publicly available at  github.com/junhongmit/KG-R1/.
@@ 1 Introduction
Retrieval-augmented generation (RAG) [gao2023retrieval] has gained popularity as a way to enhance large language models (LLMs) with access to external knowledge, thereby reducing hallucination and improving accuracy in knowledge-intensive tasks. Recent research has extended this idea to knowledge graph retrieval-augmented generation (KG-RAG), where knowledge graphs (KGs) are leveraged as the retrieval source. A KG is a structured representation of knowledge in the form of entities (nodes) and their relationships (edges) that encodes factual knowledge in a graph format. Augmenting LLMs with KGs has proven effective not only in mitigating the knowledge bottleneck but also in improving reasoning performance over complex multi-hop relations and enhancing adaptation to continually evolving real-world information [23, 16, 28]. These properties make KG-RAG especially promising in high-stakes domains, such as medical consultation and legal analysis [33, 6].
@@
As shown in Figure˜1, a typical KG-RAG adopts a modular workflow consisting of four subtasks: retrieval to query facts from KGs, reasoning to process retrieved information, reviewing to verify logical consistency, and responding to synthesize the answer [1, 23, 16, 28]. Each subtask is handled by specialized LLM modules through two main methods: (i) prompt-based modules with task-specific instructions, often including in-context demonstrations [1]; and (ii) fine-tuned modules tailored to specific tasks (e.g., SPARQL generation [7, 15] or relation extraction [30]) on specific KGs.
Figure 1: Overview of KG-R1, a multi-turn agentic KG-RAG framework. The framework enables cost-efficient inference and demonstrates strong cross-KG transferability.
@@
Despite improved reasoning accuracy, the real-world deployment of such workflows faces two key challenges: achieving high reasoning effectiveness while maintaining efficiency, and generalization to new or updated KGs. First, prompt-based methods that repeatedly call large foundation LLMs accumulate inference across stages and drive up latency, token usage, and compute (e.g., ToG [23] and ReKnoS [28]; see left panel of Figure˜2). Second, these prompted or fine-tuned modules are typically tuned to a particular KG’s domain and schema (entity types, relations, constraints), often via curated in-context examples or KG-specific fine-tuning. As a result, performance does not transfer reliably when the domain shifts, the schema changes, or the system is deployed on a new KG (e.g., RoG [16]; see right panel of Figure˜2).
Figure 2: Prior multi-module methods are costly and do not transfer well across KGs. Left: mean end-to-end generated tokens per query on WebQSP [31]. Right: average Hits@1 over five out-of-training KGQA datasets (see Section˜4.4). KG-R1 achieves both low token cost and strong cross-KG transferability.
@@
To tackle these challenges, we introduce KG-R1, an agentic KG-RAG system that employs end-to-end multi-turn reinforcement learning [12, 32, 8] to learn effective and transferable KG interaction policies. As shown in Figure˜1, the architecture of KG-R1 has two components: a single LLM agent and a KG query server. The KG query server hosts the knowledge graph along with a set of retrieval actions. The LLM agent iteratively performs cycles of short reasoning followed by retrieval actions over multiple turns, with each decision informed by knowledge obtained from the KG retrieval server, and generates a final answer. By learning a unified interaction policy, KG-R1 avoids redundant multi-module coordination and allocates computation adaptively across reasoning steps, leading to both improved effectiveness and reduced token usage. Figure˜2 demonstrates that KG-R1 achieves both high efficiency and strong cross-KG transferability simultaneously using a 3B-parameter model, outperforming prior methods. Our key contributions are summarized as follows:
@@
1. KG-R1 Framework. We introduce an agentic KG-RAG system (Section 3) that replaces multi-module pipelines with a single agent for KG-RAG, running on a lightweight KG server. The agent alternates between reasoning and retrieval actions over multiple turns, with the end-to-end trajectory optimized by RL using both turn-wise and outcome-based reward signals. Turn-wise rewards evaluate individual action effectiveness and adherence to formatting, while global rewards measure answer quality and retrieval relevance.
@@
2. Effective and Efficient Inference. By consolidating reasoning and retrieval into a single-agent, near-continuous workflow, KG-R1 achieves competitive reasoning accuracy with a small-parameter model while reducing token usage. This lowers latency and computational cost, making deployment feasible under tight budgets. Experiments demonstrate improved performance and efficiency compared to traditional multi-module workflows (Section 4.3).
@@
3. Plug-and-play Transferability across Diverse KGs. KG-R1 can be easily transferred to diverse KGs and maintains strong KG-RAG performance (Section 4.4). The trained KG-R1 agent generalizes to new KGs without retraining: backend KGs can be swapped without changing prompts, hyperparameters, or fine-tuning. This generalizability enables zero-shot transfer to unseen KGs.
@@ 2 Related Work
Knowledge Graph Retrieval-Augmented Generation (KG-RAG). KG-RAG augments LLMs with structured KGs to improve quality and compositional reasoning. Early work grounds LLMs’ generation by translating natural-language questions into executable graph queries (e.g., SPARQL/Cypher), retrieving relevant subgraphs or answers, and feeding them back to the model [18, 15]. More recent approaches adopt a modular LLM pipeline over KGs, interleaving natural language reasoning with multi-stage planning, path search, and execution, where each stage uses a prompted or fine-tuned LLM [16, 23, 4, 28]. Despite these advances, most systems rely on hand-engineered modules or prompt heuristics tied to a specific KG schema, which induce cost inefficiency and limit generalization. These challenges motivate our KG-R1 framework: a single-agent KG-RAG approach that improves efficiency and transfers well to new KGs.
Multi-turn RL for LLMs. Reinforcement learning (RL) has become central to equipping LLMs with step-by-step (chain-of-thought) reasoning behavior [17, 8]. RL-enhanced models yield substantial gains in math and coding [14, 5], and broader complex reasoning tasks. More recently, RL has been applied to agentic LLMs that invoke external tools (e.g., bash terminals and APIs) or interact with knowledge bases, improving tool use [20] and retrieval-augmented generation (RAG) [12, 26] to facilitate tool use or RAG. Building on these advances, our work, KG-R1 adopts end-to-end RL as the core optimization algorithm for an agentic KG-RAG framework.
@@ 3.1 Problem Definition
KG-RAG spans many applications, including conversational assistants [3], recommendation systems [27], and open-domain QA [35]. In this work, we instantiate and evaluate our approach on Knowledge Graph Question Answering (KGQA), which provides a grounded testbed for KG-RAG: ground-truth answers are tied to a fixed KG, evaluation is verifiable, and intermediate graph reasoning is interpretable. We now formalize knowledge graphs and KGQA tasks.
@@
Knowledge Graph Question Answering (KGQA) is a reasoning task based on KGs. Consider a dataset $D=\{(q,G,A_{q})\}$, where $q$ is a natural language question, $G$ is a KG, and $A_{q}\subseteq\mathcal{E}$ is the ground-truth answer set paired with the question. For each ground truth answer $e^{*}\in A_{q}$, there exists one or more reasoning paths in $G$ that connect anchor entities mentioned in $q$ to $e^{*}$. A reasoning path is a sequence of labeled edges $r_{1},\dots,r_{\ell}\in\mathcal{R}$ instantiated by entities $e_{0},\dots,e_{\ell}$ such that $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},$ where $e_{0}$ is an anchor entity mentioned in $q$ and $e_{\ell}=e^{*}\in A_{q}$, and $\mathcal{Z}(q,G,A_{q})$ denotes the set of all valid reasoning paths for question $q$ over $G$ with respect to ground-truth answers $A_{q}$. In solving KGQA, given $q$ and $G$, a model attempts to discover a subset of valid reasoning paths from $\mathcal{Z}(q,G,A_{q})$ and predicts $\hat{A}_{q}$ based on the terminal entities of the discovered paths. The model’s performance is evaluated by comparing $\hat{A}_{q}$ with the ground truth $A_{q}$. As an example, to answer the question “What is the capital of the U.S. state whose largest city is Chicago?”, a valid reasoning path is
$\text{Chicago}\xrightarrow{\texttt{located\_in\_state}}\text{Illinois}\xrightarrow{\texttt{capital\_of}}\text{Springfield}.$
@@ 3.2 KG-R1 Framework
Figure 3: KG-R1 framework: a single LLM agent undergoes a multi-turn generation-execution loop with a schema-agnostic KG retrieval server and responds with the final answer.
KG-R1 casts KG-RAG as a multi-turn interaction with a KG interface (KG query server). We prioritize two design principles. First, we design a single-agent architecture that simplifies deployment and enables efficient, low-cost inference. Second, we build a schema-agnostic KG retrieval server that avoids KG-specific assumptions and remains portable across different KGs. Given a KGQA dataset $D=\{(q,G,A_{q})\}$, we set up a KG retrieval server and a KG-R1 agent.
KG Query Server. The server hosts the knowledge graph $G$ and provides a set of retrieval actions $a\in\mathcal{U}_{\text{ret}}$ that enable graph traversal through 1-hop operations:
@@
The retrieval action set $\mathcal{U}_{\text{ret}}$ is sufficient for realizing any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ as a finite action sequence that arrives at $e_{\ell}\in A_{q}$ (i.e., the terminal entity is in the answer set). Forward traversal along $Z$ is implemented by $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ for $i=1,\dots,\ell$ and backward traversal is implemented by $\mathtt{get\_head\_entities}(r_{i},e_{i})$. This action design guarantees completeness of path realization, as formalized in Remark 3, and provides a schema-agnostic interaction interface that enables transfer across KGs. Full formal statements and additional properties are provided in Appendix B.
{restatable}
[Retrieval Action Set Completeness]remarkcompleteness For any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ in $G$, there exists an action sequence in $\mathcal{U}_{\text{ret}}$ of length at most $\ell{+}1$ whose output includes $e_{\ell}$.
KG-R1 Agent. We model a single LLM agent that interacts with a KG query server in a multi-turn setting. The agent undergoes initialization followed by a loop of generation and execution (see Figure 3). At initialization, a lightweight base LLM is configured with an instruction prompt $p$ (see the box below for the prompt template) containing general reasoning instructions, the user question $q$, and the KG retrieval server instructions (Table E.2). At each turn $t\leq H$, where $H$ is the maximum turn limit, the agent first undergoes the generation phase where it produces a response $m_{t}$ comprising two parts: (1) an internal reasoning wrapped in <think>...</think>, and (2) an action, which is either a KG retrieval action wrapped in <kg-query>...</kg-query>, or a final answer wrapped in <answer>...</answer>.
@@
In the following execution phase, we parse $m_{t}$ into an action $a_{t}\in\mathcal{U}_{\text{ret}}\cup\{\operatorname{Answer}\}$ using an exact-match parser $\Psi$ (i.e., $a_{t}=\Psi(m_{t})$). If $a_{t}\in\mathcal{U}_{\text{ret}}$, executing it yields an observation $o_{t+1}$ (i.e., retrieved entities, relations, or an error message in case the action does not generate a valid retrieval), which is appended to the dialogue context prior to the next turn. If $a_{t}=\operatorname{Answer}$, the interaction terminates: the content inside <answer>...</answer> is extracted and post-processed (normalization, deduplication, and entity resolution) to produce the predicted answer set $\hat{A}_{q}$. Given an interaction trajectory $\tau=(p,\,(m_{1},o_{1}),\ldots,(m_{t},o_{t}))$ with $t\leq H$, the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ is defined over token sequences $m_{t}$ and governs how textual responses are generated.